Find the equivalent resistivity of the combination of two wires of same size and length but resistivity rho1 &rho2 are connected in series
 

Grade:11

1 Answers

Arun
5 years ago
 
Dear Vinay
 
 

R1= ρ1(l/A)

R2= ρ2(l/A)

R= ρ(2l)/A

and Rs = R1+R2

» ρs = (ρ12)/2

 

Regards

Arun (askIITians forum expert)